OpenStudy (anonymous):

I don't understand how to do these problems. what is the solution set of the equation x - 11/x+4 = 2 - 4/x a. {-2, 8} b. {-8, 2} c. {-16, 1} d. {-1, 16} 2x - 4/x-2 = -2x/3 a. {3, -2} b. {2} c. {-3, 2} d. {-3}

OpenStudy (dtan5457):

Easiest way to solve these would to be plugging in those solutions and see which ones work

OpenStudy (mathmate):

I notice that the first question has only x in the equation, and all four answer have an ordered pair. Is something wrong, or perhaps I need new glasses?

OpenStudy (anonymous):

thats the equations I have

OpenStudy (dtan5457):

Yah, i've noticed that too. maybe there are 2 x values. not sure.

OpenStudy (mertsj):

\[\frac{x-11}{x+4}=2-\frac{4}{x}\] \[\frac{x-11}{x+4}=\frac{2x-4}{x}\]

OpenStudy (mertsj):

When you cross multiply you will get a quadratic equation. As a previous poster suggested, it would be easiest to just substitute the given answers into the equation and see which choice works.